# Causal Learning in Biomedical Applications

### v1

@@ Abstract
We present a benchmark for methods in causal learning. Specifically, we consider training a rich class of causal models from time-series data, and we suggest the use of the Krebs cycle and models of metabolism more broadly.
@@ 1 Introduction
Understanding causal models is important in a number of fields, from healthcare to economics, as it allows for precise forecasting a training of reinforcement learning. Learning causal models involves extracting potential non-linear relationships and dependencies between variables from sampled time-series. For example, modelling of biomarkers of non-communicable disase as a function of diet and action monitoring has shown the potential of being a powerful tool to guide the recommendations on healthy diet.
We aim to learn causal models that extend basic causal models in several directions, which are relevant in biomedical appliations. In particular, such causal models need to handle nonlinear causality, such as the competition among bacterial strains for metabolites. Likewise, one may consider constraints that become active beyond certain thresholds. Dealing with latent variables, such as hormone concentrations, is a challenge that demands a balance between model complexity and interpretability. These models should also consider cyclic relationships, time-series dynamics, and mixture-model aspects, accommodating individual variations in metabolism without predefined subgroups.
@@ 2 Preliminaries and Related Work
Throughout causal modelling, one wishes to learn a function $f$, which is known as the structural assignment map, and is closely related to the $f$ above. Under the assumption that the structural assignments are linear, noises $N_{j},j=1,\dots,N$ are indepenedntly identically distributed (i.i.d.) and follow the same Gaussian distribution, or alternatively, noises $N_{j},j=1,\dots,N$ are jointly independent, non-Gaussian with strictly positive density, one obtains linear additive noise models (ANM). In studying ANM, one may benefit from a long tradition of work on linear system identification. In particular, the identifiability of linear ANM can be reduced to identifiability of linear dynamical systems (cf. Proposition 7.5 & Theorem 7.6 in [25]).
@@
Another classic example in causal learning are Bayesian networks, first introduced by Pearl in 1985 [23]. Bayesian networks are formed by a directed acyclic graph (DAG), where each vertex $j$ represents a variable $X_{j}$, with edges going from one variable to another represent causal relationships. It is assumed that each variable $X_{j}$ is independent of other variables but for its parents ${PA}_{j}$ in the DAG, thus allowing a compressed representation of the joint probability as
@@
Instrumental variables are, however, concepts that can be used well beyond linear regression. Non-linear [20] and non-smooth [3] modifications exist. Sometimes, there is a requirement that instrumental variables might have common cofounders. This multilevel modeling is implemented in the instrumental variable toolkit by [13]. Similarly, [8] allow for a latent (hidden) variable.
@@
The tractable probabilistic models (TPMs) are a large group of methods that can be used to model probabilistic distributions compactly, in the spirit of neural networks approximating functions. A prime example of TPMs are sum-product networks (SPNs) [26], which represent the probability distribution as a DAG, where “input” random variables are assigned to leaves. Each non-leaf node corresponds to one of two operations, either sum or product. The weights of the edges are then used to learn the probability distribution. The original paper [26] also proposed an algorithm to learn the structure using a backpropagation and expectation maximization. The SPNs are only a subgroup in the broad class of probabilistic circuits [5]. The unified formalism allows to use different type of nodes besides the sum and product nodes. Dynamic versions [15, e.g.] are able to work with temporal data.
@@ 3 The Challenge
As suggested in the Introduction, we would like to learn causal models that are more expressive than many traditional models. In our view, expressivity of the causal model entails:
@@
Ability to simulate from the model entails:
@@
  • time required to simulate from the model scaling modestly (with the the number of random variables and numbers of samples).
Ability to learn the model entails:
@@
Casual Bayesian networks  [^1] Under some conditions the Bayesian networks can be used to model cyclic relationships. For example, in the case of time series, Dynamic Bayesian Networks can be used to model cyclic relationships between variables at different times. For example $X_{t+1}$ may depend on $Y_{t+1}$, which depend on $X_{t}$.
@@
In many cases, the possibility of having non-linear models is part of extensions of the original methods. A prominent example of such a method is Granger’s causality. The original method was developed with linear dependencies between the features. But further extensions were developed to include nonlinearities, for example, [32]. In Bayesian networks, the original version [23] considered only propositional variables, but subsequent versions [10, e.g.] considered also continuous variables and non-linear dependencies.
@@ 4 The Benchmark
In this paper, we present a simulated dataset based on the Krebs cycle. The Krebs cycle, also known as the citric acid cycle, is one of the fundamental pathways of biochemistry. The cycle, as illustrated in Figure 1, allows organisms that breathe to convert energy stored in food to a key energy source (ATP) in muscle cells, for example. Such a cycle presents a natural example of time series that can be used to infer causal relationships between concentrations of the reactants.
@@
Figure 1: Illustration of the Kreb’s cycle reactions used to similate the concentrations.
@@ 4.1 The Data
Depending on the modelling of the time series, each of the reactions can be represented by one or more causal relationships. Our benchmark is based on a simulator the GitHub repository at [18]. The simulator creates a virtual box with desired particles. The particles move inside the box, following the Boltzmann distribution. Once particles get close to each other, a pre-defined list of reactions is scanned to determine whether a reaction occurs, and if so, reactants are replaced with a product. The simulation then continues, and concentrations of different particles are then used as test time series. As a result, the time series contains noise (caused by the random location of particles), which is added to the locally linear behavior of the system.
In this a way, we have generated four datasets, consisting of a time series with $5$ to $5000$ time steps and $16$ features for the reactants, including $10$ in the main cycle and $6$ additional ones (incl. water). Each of the following datasets is based on simulating approximately $2500$ molecules in the bounding box:
@@
    contains $120$ series with relative concentrations, where for each triplet of the $10$ main cycle reactants, we used uniform priors, and the remaining $7$ particles were set to zero. Such a distribution is motivated by allowing the tested approaches to trace how the higher concentration of the three selected compounds moves forward in the cycle.
@@
    focuses on learning from few long time series. In this case, we have $10$ series with $5000$ time steps. We use
@@ 4.2 Evaluation Criteria
The performance of supervised machine learning methods are compared based on K-fold cross validation and F-Beta Score (or F1-score). K-Fold Cross-Validation is a technique used to assess the performance and generalizability of a machine learning model. It involves dividing the dataset into $k$ equally sized folds, training the model on $k-1$ folds, and validating it on the remaining fold. This process is repeated $k$ times, with each fold serving as the validation set once. The final performance metric is the average of the metrics obtained in each iteration.
The precision, recall, and $F_{1}$ score, which have been widely used in casual model performance measurements, are defined as
@@
$\textrm{precision}=\frac{TP}{TP+FP},$ (5)
$\textrm{recall}=\frac{TP}{TP+FN},$ (6)
where $TP$, $FP$, and $FN$ are the numbers of true positives, false positives, and false negatives, respectively.
In the case of deterministic methods, the stability of the F1-score cannot be evaluated by simple repeated evaluations followed by standard deviation calculation. Therefore, we recommend using an approach similar to cross-validation to show the stability of the results. In each evaluation, instead of plain restart, we can keep $10\,\%$ of the dataset aside to randomize data instead of the method. As a result, by doing repeated evaluations, it is possible to obtain the results’ standard deviations and confidence intervals.
@@ 4.3 A Baseline
Figure 2: An illustration of the adjacency matrix produced by various methods, and the ground truth matrix representing the set of reactions.
@@
Figure 4: An illustration of the time requriements of various methods on the Krebs dataset. The error bars show the standard deviation of the measurements calculated from $10$ repeats.
@@ 5 Conclusion
We publish all source files used to generate the data and the figures in this paper in the following GitHub repository https://github.com/petrrysavy/krebsdynotears. The repository also contains numeric results that were generated as input to the plots. The generator of the data can be found at https://github.com/petrrysavy/krebsgenerator, including a description of how to generate the benchmarking data. The generator is based on simulator at [18]. The dataset is available at https://huggingface.co/datasets/petrrysavy/krebs/tree/main.

### v2

@@ Abstract
Learning causal relationships between a set of variables is a challenging problem in computer science. Many existing artificial benchmark datasets are based on sampling from causal models and thus contain residual information that the ${R}^{2}$-sortability can identify. Here, we present a benchmark for methods in causal learning using time series. The presented dataset is not ${R}^{2}$-sortable and is based on a real-world scenario of the Krebs cycle that is used in cells to release energy. We provide four scenarios of learning, including short and long time series, and provide guidance so that testing is unified between possible users.
@@ 1 Introduction
Understanding causal models is important in a number of fields, from healthcare to economics, as it allows for precise forecasting and training of reinforcement learning algorithms. Learning causal models involves extracting potential non-linear relationships and dependencies between variables from sampled time series. For example, the modeling of biomarkers of non-communicable disease as a function of diet and action monitoring has shown the potential of being a powerful tool to guide the recommendations for a healthy diet.
Many researchers agree that there is a need for better synthetic datasets to test causal learning algorithms. This is supported by works as [30, 24, 31]. Many synthetic dataset benchmarks suffer from residual information in the data that the $R^{2}$-sortability can identify. In the case of real-world datasets, we often cannot be sure what the ground-truth causal relationships are. Often, datasets for causal discovery are too large, and as a result, they are sampled without any standardized sampling approach, thus making different papers using the datasets incomparable.
In this paper, we aim to fill this gap and provide a standardized synthetic dataset that does not suffer from the problems mentioned above. The dataset is based on simulating a set of chemical reactions describing the Krebs cycle, and for that, it uses a publicly available generator at [20]. The randomness in the data is caused by simulating the molecules in a box and providing the molecules with locations and velocities. Whenever molecules forming a left-hand side of a reaction meet, they are replaced with reactants as given by the equation.
First, we provide a brief review of a variety of methods that can be used in causal learning. Later, we provide a list of requirements that we can expect from such causal learning methods to illustrate their expressiveness. A discussion of which criteria are supported by the existing methods follows. Section 4 explains the dataset in detail and shows a possible evaluation of a method on the dataset. As an example, a state-of-the-art method named DyNoTears [25] is used. Next, we compare the presented dataset with other causal learning datasets. In conclusion, we give preference to public repositories where the dataset, as well as the source code for the evaluation of the method, can be found.
@@ 2 Preliminaries and Related Work
Throughout causal modeling, one wishes to learn a function $f$, which is known as the structural assignment map and is closely related to the $f$ above. Under the assumption that the structural assignments are linear, noises $N_{j},j=1,\dots,N$ are independently identically distributed (i.i.d.) and follow the same Gaussian distribution, or alternatively, noises $N_{j},j=1,\dots,N$ are jointly independent, non-Gaussian with strictly positive density, one obtains linear additive noise models (ANM). In studying ANM, one may benefit from a long tradition of work on linear system identification. In particular, the identifiability of linear ANM can be reduced to the identifiability of linear dynamical systems (cf. Proposition 7.5 & Theorem 7.6 in [28]).
@@
Another classic example in causal learning are Bayesian networks, first introduced by Pearl in 1985 [26]. Bayesian networks are formed by a directed acyclic graph (DAG), where each vertex $j$ represents a variable $X_{j}$, with edges going from one variable to another representing causal relationships. It is assumed that each variable $X_{j}$ is independent of other variables but for its parents, ${PA}_{j}$ in the DAG, thus allowing a compressed representation of the joint probability as
@@
Instrumental variables are, however, concepts that can be used well beyond linear regression. Non-linear [22] and non-smooth [3] modifications exist. Sometimes, there is a requirement that instrumental variables might have common cofounders. This multilevel modeling is implemented in the instrumental variable toolkit by [15]. Similarly, [9] allows for a latent (hidden) variable.
@@
The tractable probabilistic models (TPMs) are a large group of methods that can be used to model probabilistic distributions compactly in the spirit of neural networks approximating functions. A prime example of TPMs is sum-product networks (SPNs) [29], which represent the probability distribution as a DAG, where “input” random variables are assigned to leaves. Each non-leaf node corresponds to one of two operations, either sum or product. The weights of the edges are then used to learn the probability distribution. The original paper [29] also proposed an algorithm to learn the structure using backpropagation and expectation maximization. The SPNs are only a subgroup in the broad class of probabilistic circuits [6]. The unified formalism allows using different types of nodes besides the sum and product nodes. Dynamic versions [17, e.g.] are able to work with temporal data.
@@ 3 The Challenge
As suggested in the Introduction, we would like to learn causal models that are more expressive than many traditional models. In our view, the expressivity of the causal model entails:
@@
The ability to simulate from the model entails:
@@
  • time required to simulate from the model scaling modestly (with the number of random variables and numbers of samples).
The ability to learn the model entails:
@@
Causal Bayesian networks  [^1] Under some conditions the Bayesian networks can be used to model cyclic relationships. For example, in the case of time series, Dynamic Bayesian Networks can be used to model cyclic relationships between variables at different times. For example $X_{t+1}$ may depend on $Y_{t+1}$, which depend on $X_{t}$.
@@
In many cases, the possibility of having non-linear models is part of extensions of the original methods. A prominent example of such a method is Granger’s causality. The original method was developed with linear dependencies between the features. But further extensions were developed to include nonlinearities, for example, [37]. In Bayesian networks, the original version [26] considered only propositional variables, but subsequent versions [12, e.g.] considered also continuous variables and non-linear dependencies.
@@ 4 The Benchmark
In this paper, we present a simulated dataset based on the Krebs cycle. The Krebs cycle, also known as the citric acid cycle, is one of the fundamental pathways of biochemistry. The cycle, as illustrated in Figure 1, allows organisms that breathe to convert the energy stored in food to a key energy source (ATP) in muscle cells, for example. Such a cycle presents a natural example of time series that can be used to infer causal relationships between concentrations of the reactants.
@@
Figure 1: Illustration of the Kreb’s cycle reactions used to simulate the concentrations.
@@ 4.1 The Data
Depending on the modeling of the time series, each of the reactions can be represented by one or more causal relationships. Our benchmark is based on a simulator in the GitHub repository at [20]. The simulator creates a virtual box with desired particles. The particles move inside the box, following the Boltzmann distribution. Once particles get close to each other, a pre-defined list of reactions is scanned to determine whether a reaction occurs, and if so, reactants are replaced with a product. The simulation continues, and concentrations of different particles are used as test time series. As a result, the time series contains noise (caused by the random location of particles), which is added to the locally linear behavior of the system.
In this way, we have generated four datasets, consisting of a time series with $5$ to $5000$ time steps and $16$ features for the reactants, including $10$ in the main cycle and $6$ additional ones (incl. water). Each of the following datasets is based on simulating approximately $2500$ molecules in the bounding box:
@@
    contains $120$ series with relative concentrations, where for each triplet of the $10$ main cycle reactants, we used uniform priors, and the remaining $7$ particles were set to zero. Such a distribution is motivated by allowing the tested approaches to trace how the higher concentration of the three selected compounds move forward in the cycle.
@@
    focuses on learning from a few long time series. In this case, we have $10$ series with $5000$ time steps. We use
@@ 4.2 Evaluation Criteria
For comparison, the dataset includes the ground-truth causal matrix as defined by the equations. Also, please note diagonal in Fig. 2(a), which is there as well because the presence of a substance at time $t$ implies presence of the same substance at time $t+1$. A single run of an algorithm produces a causal matrix that can be compared to the ground truth one. One of the simplest measures to compare the matrices is the structural hamming distance (SDH), that counts the number of edges that need to be added and the number of edges that need to be removed to convert the predicted causal graph into the ground-truth causal graph.
We propose as the main measure of the quality of the causal matrix to use the $F1$-score, which is the harmonic mean of the precision and recall measures. Let TP, TN, FP, FN be the true/false positive/negative measures as in a classification task. Then, the $F1$-score is defined
@@
precision $=\frac{TP}{TP+FP},$ (5)
recall $=\frac{TP}{TP+FN}.$ (6)
Please note, that this measure can be easily extended to the case when the predicted causal matrix is stochastic. In that case, for example, an edge predicted with weight $0.3$ when there is no ground-truth edge, contributes $0.3$ to false-positive and $0.7$ to true-negative.
To assess the stability of the method, we recommend to average the results over at least $10$ runs of the method, whenever the tested method is randomized. The standard mean should then be calculated. In the case of deterministic methods, the stability of the $F1$-score cannot be evaluated by simple repeated evaluations followed by standard deviation calculation. Therefore, we recommend using an approach similar to cross-validation to show the stability of the results. In each evaluation, instead of plain restart, we can keep $10\,\%$ of the dataset aside to randomize data instead of the method. As a result, by doing repeated evaluations, it is possible to obtain the results’ standard deviations and confidence intervals.
@@ 4.3 A Baseline
Figure 2: An illustration of the adjacency matrix produced by various methods and the ground truth matrix representing the set of reactions. Black squares represent $1$ an edge in the adjacency matrix, grey 00.
@@
Figure 4: An illustration of the time requirements of various methods on the Krebs dataset. The error bars show the standard deviation of the measurements calculated from $10$ repeats.
5 Discussion
Once the dataset is presented, we are ready to compare it with other existing possibilities and show how it improves upon the other choices in [25, 11, 4]. We will point out the important advantages that the Krebs dataset has against other datasets.
Our method does not assume any ground truth structural model.
Instead, our method uses an independent method of simulation from a real-world setting. The dataset is generated by following the chemical reactions in the Krebs cycle. This makes it possible to generate multiple variants (KrebsN, Krebs3, KrebsS, KrebsS) consistently. These consist of a time series with $5$ to $5000$ time steps and $16$ features for the reactants, including $10$ in the main cycle and $6$ additional ones (incl. water).
Dataset $R^{2}$-sortability Standard variance Note
krebsN $0.486$ $0.008$
krebs3 $0.501$ $0.011$
krebsS $0.497$ $0.035$ (please, see caption)
krebsL $0.492$ $0.005$
Table 3: The evaluation of $R^{2}$-sortability for individual time series. For each of the time series in each dataset, we calculated the $R^{2}$-sortability using the CausalDisco Python package [30, 31]. To obtain the results, the $R^{2}$-sortability values were then averaged over each of the datasets. Please note that for 11 time series in the krebsS dataset, the $R^{2}$-sortability method did not produce a numeric result. Since this is much less than $1\,\%$ of the dataset (and $R^{2}$-sortability is bounded by 00 and $1$), the average won’t be influenced substantially.
The presented dataset is not $R^{2}$-sortable.
Our method does not suffer from the $R^{2}$-sortability issues other synthetic benchmarks suffer from, as explained by [24] and [31]. Indeed, Ormaniec et al. [24] argue that there are usually patterns left by the simulation from structural models that are easy to exploit. This can be quantified by the $R^{2}$-sortability [31]. To illustrate how the Krebs dataset stands compared to the $R^{2}$-sortability, we implemented a code evaluating the $R^{2}$-sortability for our dataset, the results of which can be seen in Table 3. Reference [31] then explains that “0.5 means that ordering the variables by $R^{2}$ amounts to a random guess of the causal ordering”, meaning that our dataset is not $R^{2}$-sortable. Thus, the fact that we do not assume any underlying framework makes our dataset more universal.
The ground truth causal relationships are known.
At the same time, our method comes with widely accepted ground-truth data. The advantage can be seen when compared to datasets such as S&P100 (stock returns for 100 top US companies), used in DyNoTears paper [25]. S&P100 is a real-world dataset that suffers from an unclear ground truth causal matrix. Moreover, the authors had to ensure that the data were stationary, as concept drift is likely to happen in stock trading.
A similar situation is connected with the SACHS dataset [4]. This dataset contains single-cell measurements of levels of 11 proteins in immune cells. With  853 samples, the dataset is of a similar size to ours. However, we cannot be sure what the true causal relationships between the variables representing individual genes are in the case of expression data.
Prospect of perfect reconstruction
At the same time, our method allows for the prospect of perfect reconstruction. Our dataset is much smaller than another commonly used causality dataset, the DREAM dataset [11]. This is desirable in connection with the fact that most of the problems in causal learning are NP-hard. Because of that, perfect recovery with many variables is computationally infeasible. Causal discovery algorithms should be tested on smaller, easy-to-explain datasets first before proceeding to larger and more complex datasets. The use of larger datasets also brings another reproducibility problem – sampling, often done in an ad hoc, paper-specific fashion – which is not needed with our data.
@@ 6 Conclusion
We publish all source files used to generate the data and the figures in this paper in the following GitHub repository https://github.com/petrrysavy/krebsdynotears. The repository also contains numeric results that were generated as input to the plots. The generator of the data can be found at https://github.com/petrrysavy/krebsgenerator, including a description of how to generate the benchmarking data. The generator is based on a simulator at [20]. The dataset is available at https://huggingface.co/datasets/petrrysavy/krebs/tree/main.